# MotionCLR: Motion Generation and Training-free Editing via Understanding Attention Mechanisms

### v1

@@ C.1 Mathematical Visualization of Self-attention Mechanism
(1) Similarity Calculation. In the first step, the similarity between the motion embeddings at different frames is computed using the dot product, represented by $\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}=\mathbf{{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}Q}{\color[rgb]{0.9765625,0.828125,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9765625,0.828125,0.703125}K}}^{\top}$. This measurement reflects the internal relationship/similarity between different motion frames within the sequence. Fig. 23(a) illustrates how the $\mathtt{softmax}(\cdot)$ operation is applied to the similarity matrix to determine which motion feature should be selected at a given frame $f$.
(2) Feature Updating. Next, the similarity scores are used to weight the motion embeddings ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$) and generate updated features $\mathbf{\color[rgb]{0.92578125,0.71484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.71484375,0.515625}X^{\prime}}$, as shown by the equation $\mathbf{\color[rgb]{0.92578125,0.71484375,0.515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.92578125,0.71484375,0.515625}X^{\prime}}=\mathtt{softmax}(\mathbf{{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}Q}{\color[rgb]{0.9765625,0.828125,0.703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.9765625,0.828125,0.703125}K}}^{\top}/\sqrt{d})\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$. Here, the similarity matrix applies its selection of values ($\mathbf{\color[rgb]{0.96484375,0.7421875,0.5390625}\definecolor[named]{pgfstrokecolor}{rgb}{0.96484375,0.7421875,0.5390625}V}$) to update the motion features. This process allows the self-attention mechanism to dynamically adjust the representation of each motion frame based on its relevance to other frames in the sequence.
C.2 Mathematical Visualization of Cross-attention Mechanism
(a) $\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}=\mathbf{{\color[rgb]{1,0.6015625,0.24609375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.24609375}Q}{\color[rgb]{0.94921875,0.62109375,0.66015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94921875,0.62109375,0.66015625}K}}^{\top}$.
(b) $\mathbf{\color[rgb]{0.32421875,0.40625,0.6953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.32421875,0.40625,0.6953125}X^{\prime}}=\mathtt{softmax}(\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}/\sqrt{d})\mathbf{{\color[rgb]{0.703125,0.8671875,0.58203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.703125,0.8671875,0.58203125}V}}$.
@@ C.2 Mathematical Visualization of Cross-attention Mechanism
(1) Similarity Calculation. In the first step, the similarity between the motion embeddings ($\mathbf{{\color[rgb]{1,0.6015625,0.24609375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.24609375}Q}}$) with $F$ frames and the text embeddings ($\mathbf{{\color[rgb]{0.94921875,0.62109375,0.66015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94921875,0.62109375,0.66015625}K}}$) with $N$ words is computed through the dot product, represented by $\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}=\mathbf{{\color[rgb]{1,0.6015625,0.24609375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.24609375}Q}{\color[rgb]{0.94921875,0.62109375,0.66015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94921875,0.62109375,0.66015625}K}}^{\top}$. This similarity measurement reflects the relationship between motion frames and words. Fig. 24(a) shows how the $\mathtt{softmax}(\cdot)$ operation is applied to the similarity matrix to determine which word should be activated at a given frame $f$.
As discussed in the main text (see Sec. 1), despite the progresses in human motion generation [Zhang et al., 2024d, Cai et al., 2024, Zhang et al., 2024c, Guo et al., 2024c, Raab et al., 2024b, Kapon et al., 2024, Cohan et al., 2024, Fan et al., 2024, Xu et al., 2024, 2023a, 2023b, Yao et al., 2022, Feng et al., 2023, Ao et al., 2023, Yao et al., 2024, Zhang et al., 2024f, Liu et al., 2010, Aberman et al., 2020a, Karunratanakul et al., 2024, Li et al., 2024, 2023a, Gong et al., 2023, Zhou and Wang, 2023, Zhong et al., 2023, Zhu et al., 2023, Athanasiou et al., 2023, Zhong et al., 2024, Guo et al., 2024b, Zhang et al., 2024c, Athanasiou et al., 2024, Zhao et al., 2023, Zhang et al., 2022, 2020, Diomataris et al., 2024, Pinyoanuntapong et al., 2024, Diller and Dai, 2024, Peng et al., 2023, Hou et al., 2023, Liu et al., 2023, Cong et al., 2024, Cui et al., 2024, Jiang et al., 2022, Kulkarni et al., 2024, Tessler et al., 2024, Liang et al., 2024, Ghosh et al., 2023, Wu et al., 2024], there still lacks a explicit modeling of word-level cross-modal correspondence in previous work. To clarify this, our method models a fine-grained word-level cross-modal correspondence.
@@ C.3 The Basic Difference with Previous Diffusion-based Motion Generation Models in Cross-modal Modeling
As illustrated in Fig. 25, the major distinction between our proposed method and previous diffusion-based motion generation models lies in the explicit modeling of word-level cross-modal correspondence. In the MDM-like fashion Tevet et al. [2022b] (see Fig. 25a), previous methods usually utilize a denoising transformer encoder that treats the entire text as a single embedding, mixing it with the motion sequence. This approach lacks the ability to capture the nuanced relationship between individual words and corresponding motion elements, resulting in an over-compressed representation. Although we witness that Zhang et al. [2024b] also introduces cross-attention in the motion generation process, it still faces two problems in restricting the fine-grained motion editing applications. First of all, the text embeddings are mixed with frame embeddings of diffusion, resulting in a loss of detailed semantic control. Our approach disentangles the diffusion timestep injection process in the convolution module to resolve this issue. Besides, the linear cross-attention in MotionDiffuse is different from the computation process of cross-attention, resulting in a lack of explanation of the word-level cross-modal correspondence. The auto-regressive (AR) fashion [Zhang et al., 2023a] (Fig. 25b) adopts a simple concatenation of text and motion, where an AR transformer processes them together. However, this fashion also fails to explicitly establish a fine-grained correspondence between text and motion, as the AR transformer merely regards the text and motion embeddings as one unified sequence.
@@ Appendix E Implementation and Evaluation Details
E.1 Implementation Details
The MotionCLR model is trained on the HumanML3D dataset with one NVIDIA A-100 GPU based on PyTorch [Paszke et al., 2019]. The latent dimension of the motion embedding is $512$. We take the CLIP-ViT-B model to encoder text as word-level embeddings. The training process utilizes a batch size of 64, with a learning rate initialized at $42e-$ 4 and decaying at a rate of $0.9$ every $5,000$ steps. Additionally, a weight decay of $21e-$ 2 is employed to regularize the model parameters. For the diffusion process, the model is trained over $1,000$ diffusion steps. We incorporate a probability of $0.1$ for condition masking to facilitate classifier-free guidance learning. During training, dropout is set at $0.1$ to prevent overfitting, and all networks in the architecture follow an $8$-layer Transformer design.
[…]
E.2 Compared Baselines
@@
TM2T [Guo et al., 2022b] explores the reciprocal generation of 3D human motions and texts. It uses motion tokens for compact representation, enabling flexible generation for both text2motion and motion2text tasks.
T2M [Guo et al., 2022a] generates diverse 3D human motions from text using a two-stage approach involving text2length sampling and text2motion generation. It employs a motion snippet code to capture semantic contexts for more faithful motion generation.
@@
MLD [Chen et al., 2023b] uses a diffusion process on motion latent space for conditional human motion generation. By employing a Variational AutoEncoder (VAE), it efficiently generates vivid motion sequences while reducing computational overhead.
MotionDiffuse [Zhang et al., 2024b] is a diffusion model-based text-driven framework for motion generation. It provides diverse and fine-grained human motions, supporting probabilistic mapping and multi-level manipulation based on text prompts.
T2M-GPT [Zhang et al., 2023a] combines a VQ-VAE and GPT to generate human motions from the text. With its simple yet effective design, it achieves competitive performance and outperforms some diffusion-based methods on specific metrics.
@@
E.3 Evaluation Details
Motion (de-)emphasizing. To evaluate the effectiveness of motion (de-)emphasizing application, we construct 100 prompts to verify the algorithm. All of these prompts are constructed by researchers manually. We take some samples from our evaluation set as follows.
@@ E.3 Evaluation Details
Each line in the examples represents the index of the edited word in the sentence, followed by the corresponding prompt. These indices indicate the key verbs or actions that are subject to the (de-)emphasizing during the evaluation process. The prompts were carefully selected to cover a diverse range of actions, ensuring that our method is tested on different types of motion descriptions. For instance, in the prompt “3 the figure leaps high”, the number 3 indicates that the word “leaps” is the third word in the sentence and is the target action for (de-)emphasizing. This format ensures a systematic evaluation of how the model responds to adjusting attention weights on specific actions across different prompts.
@@
Action counting. To thoroughly evaluate the effectiveness of our action counting method, we constructed a test set containing 70 prompts. These prompts were manually designed by researchers to ensure diversity. Each prompt corresponds to a motion sequence generated by our model, and the ground truth action counts were labeled by researchers based on the observable actions within the generated motions.
@@ F.2 Supplementary for Motion Style Transfer
As discussed in the main text, motion style transfer is accomplished by replacing the query ($\mathbf{Q}$) from the content sequence ($\mathbf{M_{2}}$) with that from the style sequence ($\mathbf{M_{1}}$). This replacement ensures that while the content features from $\mathbf{M_{2}}$ are preserved, the style features from $\mathbf{M_{1}}$ are adopted, resulting in a synthesized motion sequence that captures the style of $\mathbf{M_{1}}$ with the content of $\mathbf{M_{2}}$.
Fig. 28 provides a visual explanation of this process. The self-attention mechanism plays a crucial role, where the attention map determines the correspondence between the style and content features. The pseudo code snippet provided in Code LABEL:lst:style_transfer exemplifies this process. By setting “query[1] = query[0]” in the code, the query for the second motion ($\mathbf{M_{2}}$) is replaced by that of the first motion ($\mathbf{M_{1}}$), which effectively transfers the motion style from $\mathbf{M_{2}}$ to $\mathbf{M_{1}}$. In summary, this motion style transfer method allows one motion sequence to adopt the style characteristics of another while maintaining its own content.

### v2

@@ C.3 The Basic Difference with Previous Diffusion-based Motion Generation Models in Cross-modal Modeling
(c) $\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}=\mathbf{{\color[rgb]{1,0.6015625,0.24609375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.6015625,0.24609375}Q}{\color[rgb]{0.94921875,0.62109375,0.66015625}\definecolor[named]{pgfstrokecolor}{rgb}{0.94921875,0.62109375,0.66015625}K}}^{\top}$.
(d) $\mathbf{\color[rgb]{0.32421875,0.40625,0.6953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.32421875,0.40625,0.6953125}X^{\prime}}=\mathtt{softmax}(\mathbf{\color[rgb]{0.6875,0.6875,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.6875,0.6875,0.6875}S}/\sqrt{d})\mathbf{{\color[rgb]{0.703125,0.8671875,0.58203125}\definecolor[named]{pgfstrokecolor}{rgb}{0.703125,0.8671875,0.58203125}V}}$.
@@ Appendix E Implementation and Evaluation Details
E.1 Compared Baselines
@@
MLD [Chen et al., 2023] uses a diffusion process on motion latent space for conditional human motion generation. By employing a Variational AutoEncoder (VAE), it efficiently generates vivid motion sequences while reducing computational overhead.
MotionDiffuse [Zhang et al., 2024a] is a diffusion model-based text-driven framework for motion generation. It provides diverse and fine-grained human motions, supporting probabilistic mapping and multi-level manipulation based on text prompts.
@@
E.2 Evaluation Details
[…]
@@ E.2 Evaluation Details
Each line in the examples represents the index of the edited word in the sentence, followed by the corresponding prompt. These indices indicate the key verbs or actions that are subject to the (de-)emphasis during the evaluation process. The prompts were carefully selected to cover a diverse range of actions, ensuring that our method is tested on different types of motion descriptions. For instance, in the prompt “3 the figure leaps high”, the number 3 indicates that the word “leaps” is the third word in the sentence and is the target action for (de-)emphasis. This format ensures a systematic evaluation of how the model responds to adjusting attention weights on specific actions across different prompts.
@@
Action counting. To thoroughly evaluate the effectiveness of our action counting method, we constructed a test set based on HVerb-wild. These prompts were manually designed by researchers to ensure diversity. Each prompt corresponds to a motion sequence generated by our model, and the ground truth action counts were labeled by researchers based on the observable actions within the generated motions.
E.3 User Study Details
To evaluate the effectiveness of our editing methods, we conducted a user study with participants rating the quality of edited motion results. Fig. 29 illustrates an example of the interface and questions used during the study (replacement case for example here). Participants were asked to compare the original motion (leftmost column) with the results of 3 editing methods and rate them across three metrics (mentioned in the main text). This study provides insight into the subjective perception of motion quality and highlights the differences between various editing approaches.
Figure 29: An example interface from the user study.
@@ F.2 Supplementary for Motion Style Transfer
As discussed in the main text, motion style transfer is accomplished by replacing the query ($\mathbf{Q}$) from the content sequence ($\mathbf{M_{2}}$) with that from the style sequence ($\mathbf{M_{1}}$). This replacement ensures that while the content features from $\mathbf{M_{2}}$ are preserved, the style features from $\mathbf{M_{1}}$ are adopted, resulting in a synthesized motion sequence that captures the style of $\mathbf{M_{1}}$ with the content of $\mathbf{M_{2}}$. Fig. 30 provides a visual explanation of this process. The self-attention mechanism plays a crucial role, where the attention map determines the correspondence between the style and content features. The pseudo code snippet provided in Code LABEL:lst:style_transfer exemplifies this process. By setting “query[1] = query[0]” in the code, the query for the second motion ($\mathbf{M_{2}}$) is replaced by that of the first motion ($\mathbf{M_{1}}$), which effectively transfers the motion style from $\mathbf{M_{2}}$ to $\mathbf{M_{1}}$. In summary, this motion style transfer method allows one motion sequence to adopt the style characteristics of another while maintaining its content.
@@ Appendix G Details of Action Counting in a Motion
Appendix H Web User Interface for Interactive Motion Generation and Editing
To have a better understanding of our task, we build a user interface with Gradio [Abid et al., 2019]. We introduce the demo as follows. In Fig. 31, we illustrate the steps involved in generating and visualizing motions using the interactive interface. Fig. 31(a) displays the initial step where the user provides input text such as “a man jumps” and adjusts motion parameters. Once the settings are finalized, the system begins processing the motion based on these inputs, as seen in the left panel. Fig. 31(b) showcases the generated motion based on the user’s input. The interface provides a rendered output of the skeleton performing the described motion. This presentation allows users to easily correlate the input parameters with the resulting animation. The generated motion can further be edited by adjusting parameters such as the length of the motion, emphasizing or de-emphasizing certain actions, or replacing actions altogether, depending on user requirements. This process demonstrates how the interface facilitates a workflow from input to motion visualization.
(a) Motion generation interface example.
(b) Generated Motion Example
Figure 31: Motion generation and its output examples.
The logical sequence of operations is as follows:
  1. 1.
  2. 2.
  3. 3.
    Motion editing: We show some downstream tasks of MotionCLR here.
(a) Motion (de-)Emphasizing interface.
(b) In-place replacement example.
(c) Example-based motion generation progress.
(d) Example-based motion generation results.